# NPU-NTU System for Voice Privacy 2024 Challenge

### v1

@@ 3.4 Anonymization Strategy
Table 1: Averaged results over baseline systems and our proposed anonymization system on VPC 2024 development and test datasets. F and M represent the female and male, respectively. Avg denotes average results between female and male EER results. B* denotes different baseline systems described in VPC 2024 and C* represents the different EER conditions.
@@ 4.2 Metrics
The equal error rate (EER) is used as the privacy metric, and the word error rate (WER) and unweighted average recall (UAR) are used as utility metrics. A higher EER indicates better privacy protection, a lower WER indicates better intelligibility, and a higher UAR indicates better emotion preservation. All evaluation setups follow the same configuration as the VPC 2024 pipelines [7].

### v2

@@ 3.4 Anonymization Strategy
Table 1: Averaged results over baseline systems and our proposed anonymization system on VPC 2024 development and test datasets. F and M represent the female and male, respectively. Avg denotes average results between female and male EER results and B* denotes different baseline systems described in VPC 2024. C* represents the different EER conditions under VPC 2024 setting and S* represents the different EER conditions under the semi-informed scenario.
@@ 4.2 Metrics
The equal error rate (EER) is used as the privacy metric, and the word error rate (WER) and unweighted average recall (UAR) are used as utility metrics. A higher EER indicates better privacy protection, a lower WER indicates better intelligibility, and a higher UAR indicates better emotion preservation. Furthermore, we employ two types of privacy protection scenarios: semi-informed and white box. These settings follow VPC 2022 [6] and VPC 2024 [7] respectively. In VPC 2024, attackers know all hyper-parameters of the anonymization system, which is a challenging scenario.